Scottsdale Ltd. manufactures a variety of high-volume and low-volume products to customer demand. Presented is information on 2009 manufacturing overhead and activity cost drivers.

Level

Total Cost

Units of Cost Driver

Unit

$600,000

20,000 machine hours

Batch

200,000

1,000 customer orders

Product

400,000

50 products

Product X1 required 3,000 machine hours to fill 10 customer orders for a total of 8,000 units.

(a) Assuming all manufacturing overhead is estimated and predicted on the basis of machine hours, determine the predicted total overhead costs to produce the 8,000 units of product X1.
$Answer

(b) Assuming manufacturing overhead is estimated and predicted using separate rates for machine hours, customer orders, and products (a multiple-level cost hierarchy), determine the predicted total overhead costs to produce the 8,000 units of product X1.
$Answer

(c) Calculate the error in predicting manufacturing overhead using machine hours versus using multiple cost drivers. Indicate whether the use of only machine hours results in overpredicting or underpredicting the costs to produce 8,000 units of product X1.
$Answer AnswerOver-predictionUnder-prediction

(d) Determine the error in the prediction of X1 batch-level costs resulting from the use of only machine hours. Indicate whether the use of only machine hours results in overpredicting or underpredicting the batch-level costs of product X1.
$Answer AnswerOver-predictionUnder-prediction

(e) Determine the error in the prediction of X1 product-level costs resulting from the use of only machine hours. Indicate whether the use of only machine hours results in overpredicting or underpredicting the product-level costs of product X1.
$Answer AnswerOver-predictionUnder-prediction
Answer #1

(a) Predetermined overhead rate = Total manufacturing ocerhead cost/Total machine-hours = ($600000 + $200000 + $400000)/20000 = $1200000/20000 = $60 per machine-hour

Predicted total overhead costs to produce 8000 units of X1 = 3000 X $60 = $180000

(b) Predicted total overhead costs to produce 8000 units of X1 = $100000

Level Total Cost Units of Cost Driver Activity Rate Product X1
Cost Driver Usage Overhead Costs
Unit 600000 20000 30 3000 90000
Batch 200000 1000 200 10 2000
Product 400000 50 8000 1 8000
Total 100000

(c) $180000 - $100000 = $80000 Over-prediction

(d) Batch-level: ($200000 x 3000/20000) - $2000 = $30000 - $2000 = $28000 Over-prediction

(e) Product-level: ($400000 x 3000/20000) - $8000 = $60000 - $8000 = $52000 Over-prediction